### v1

# Model Predictive Control via Probabilistic Inference: A Tutorial

Abstract
Model Predictive Control (MPC) is a fundamental framework for optimizing robot behavior over a finite future horizon. While conventional numerical optimization methods can efficiently handle simple dynamics and cost structures, they often become intractable for the nonlinear or non-differentiable systems commonly encountered in robotics. This article provides a tutorial on probabilistic inference-based MPC, presenting a unified theoretical foundation and a comprehensive overview of representative methods. Probabilistic inference-based MPC approaches, such as Model Predictive Path Integral (MPPI) control, have gained significant attention by reinterpreting optimal control as a problem of probabilistic inference. Rather than relying on gradient-based numerical optimization, these methods estimate optimal control distributions through sampling-based techniques, accommodating arbitrary cost functions and dynamics. We first derive the optimal control distribution from the standard optimal control problem, elucidating its probabilistic interpretation and key characteristics. The widely used MPPI algorithm is then derived as a practical example, followed by discussions on prior and variational distribution design, tuning principles, and theoretical aspects. This article aims to serve as a systematic guide for researchers and practitioners seeking to understand, implement, and extend these methods in robotics and beyond.
keywords:
Model Predictive Control, Probabilistic Inference, Sampling-Based MPC, Model Predictive Path Integral Control, Model-Based Reinforcement Learning
1 Introduction
Establishing a universal control theory applicable to diverse robots and tasks has long been a fundamental goal in robotics research. For complex tasks, we often require control methods that optimize actions while predicting future states. Model Predictive Control (MPC) and related optimal control frameworks are built upon this principle. MPC formulates robot control and task execution as finite-horizon optimal control problems, solving them at each time step to determine sophisticated robot behaviors.
To apply MPC to a specific application, we design cost functions and system dynamics for the target robot and task. However, the resulting formulations often exhibit strong nonlinearity, making them intractable for numerical optimization methods. In many practical applications, what we truly need is not infinite-horizon stability or convergence guarantees, but rather an optimization framework capable of reliably solving a broad class of optimal control problems with straightforward implementation.
This article provides a tutorial on the theory and applications of probabilistic inference-based approaches to solving MPC optimal control problems, a topic that has attracted robotic research attention in recent years. This framework belongs to the class of sampling-based MPC methods, which randomly sample robot actions and determine optimal behaviors from these samples. Consequently, it is fundamentally applicable to optimal control problems with arbitrary dynamics and cost functions.
Attempts to solve optimal control problems through probabilistic inference date back to the early 2000s, beginning with path integral control [kappen2005path], extending to information-theoretic frameworks through Model Predictive Path Integral Control (MPPI) [williams2018information], and reaching connections to Reinforcement Learning (RL) through the general framework of Control as Inference (CaI) [levine2018reinforcement]. This article begins with an exposition of general frameworks such as CaI, derives the widely-used MPPI algorithm, and introduces advanced methods and practical considerations for applications. In addition to the aforementioned references, this work draws upon [wang2021variational, abdolmaleki2018maximum].
2 Preliminaries
Figure 1: Overview of probabilistic inference-based MPC. The framework consists of two stages: (1) deriving the optimal control distribution from the optimal control problem, and (2) generating control input sequences from the optimal control distribution.
2.1 Optimal Control Problem Formulation
Our objective is to find the optimal control input (action) sequence $\mathbf{u}^{*}_{0:T-1}$ by solving the following optimal control problem over a finite time horizon $T$:
$\min_{\mathbf{u}_{0:T-1}}\;$ $J\left(\mathbf{x}_{0:T},\mathbf{u}_{0:T-1}\right),$ (1a)
s.t. $\mathbf{x}_{t+1}=\mathbf{f}(\mathbf{x}_{t},\mathbf{u}_{t}),\quad^{\forall}t\in\{0,\dots,T-1\}.$ (1b)
Here, $J$ denotes the cost function, $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ represent the state and control input at time $t$, $\mathbf{u}_{0:T-1}=[\mathbf{u}_{0},\dots,\mathbf{u}_{T-1}]$ is the control input sequence, $\mathbf{x}_{0:T}=[\mathbf{x}_{0},\dots,\mathbf{x}_{T}]$ is the state sequence, $\mathbf{f}$ represents the system dynamics, and $\mathbf{x}_{0}$ is the observable initial state. Note that most optimal control problems include constraints on states and control inputs. In probabilistic inference MPC, input constraints can be directly incorporated as part of the dynamics, while state constraints cannot be handled directly and are typically embedded in the cost function using indicator functions. For details, refer to the experiments in [williams2018information].
Users design the cost function $J$ and dynamics $\mathbf{f}$, solve Eq. (1) in real-time, and control the robot by applying $\mathbf{u}^{*}_{0:T-1}$. When $J$ and $\mathbf{f}$ have simple forms (e.g., quadratic), efficient solutions can be obtained using numerical optimization methods. However, when $J$ or $\mathbf{f}$ exhibit nonlinearity or lack analytical derivatives, solving via numerical optimization becomes extremely challenging.
2.2 Motivations for Probabilistic Inference-Based MPC
When numerical optimization proves infeasible, our remaining option is computationally intensive sampling-based methods. The most straightforward approach is random shooting method, i.e., randomly generate a large number of control input samples $\mathbf{u}_{0:T-1}$, simulate state trajectories using Eq. (1b), compute the cost $J$ for each trajectory via Eq. (1a), and select the control input sequence with the minimum cost. However, as shown in Fig. 2, this method suffers from extremely poor sample efficiency and faces the curse of dimensionality, where the required number of samples grows exponentially with input dimensionality.
Figure 2: A comparison of sampling-based MPC methods in a vehicle obstacle avoidance task [honda2024stein]. Left: Random shooting method. Right: Model Predictive Path Integral control (MPPI). Random shooting method finds a single optimal solution from random samples, leading to poor sample efficiency. In contrast, MPPI identifies an optimal control distribution, improving sample efficiency significantly.
To overcome the sample efficiency challenges, We therefore introduce a probabilistic inference framework. Specifically, as illustrated in Fig. 1, we approximate the probabilistic distribution of optimal control inputs (hereafter referred to as the optimal control distribution) using samples and generate control input sequences from this distribution. Instead of seeking a single optimal solution from random samples, identifying a distribution with inherent breadth based on samples significantly improves sample efficiency, as shown in Fig. 2. Beyond improved sample efficiency, formulating the problem as distribution optimization offers several additional advantages:
  • Enables representation of stochastic behavioral variability.
  • Consists primarily of parallel independent operations, facilitating straightforward parallelization on CPUs/GPUs.
  • Allows end-to-end differentiable implementation using libraries such as PyTorch.
In particular, stochastic behavioral variability is valuable for representing exploration behavior and plays a crucial role in model-based reinforcement learning [hansen2022temporal, okada2020variational].
3 Probabilistic Inference-Based MPC
In this section, we introduce the foundations of the probabilistic inference-based MPC framework. As shown in Fig. 1, the scheme involves two key steps: deriving the optimal control distribution from the optimal control problem in Eq. (1) and then generating control inputs to apply to the robot from this optimal control distribution. Thus, two fundamental questions arise:
  1. 1.
    How is the optimal control distribution represented, and what are its characteristics?
  2. 2.
    How do we generate control inputs to apply to the robot from the optimal control distribution?
The following exposition focuses on these two fundamental questions.
3.1 Derivation of the Optimal Control Distribution
We begin by deriving the optimal control distribution from the optimal control problem in Eq. (1). First, we state the conclusion: under some assumptions, the probability density function $\pi^{*}$ of the optimal control distribution can be expressed by the following remarkably simple formula:
$\pi^{*}(\mathbf{u}_{0:T-1})=Z^{-1}\underbrace{\exp\left(-\lambda^{-1}J_{\tau}(\mathbf{u}_{0:T-1})\right)}_{\text{Boltzmann distribution}}\underbrace{p(\mathbf{u}_{0:T-1})}_{\text{Prior distribution}}.$ (2)
Here, $Z$ is the normalization constant, $\lambda$ is the temperature parameter, $J_{\tau}(\mathbf{u}_{0:T-1})$ is the trajectory cost obtained from the control input $\mathbf{u}_{0:T-1}$ via Eq. (1), and $p(\mathbf{u}_{0:T-1})$ represents the prior distribution over control inputs. Thus, the optimal control distribution is expressed as the product of a Boltzmann distribution and a prior distribution [^1] A similar expression appears in Direct Preference Optimization [rafailov2023direct], known for alignment methods in large language models.. While we detail the Boltzmann distribution later, it is a distribution that expresses trajectory optimality. We now derive this expression in three steps.
3.1.1 Graphical Model Representation of the Optimal Control Problem
Figure 3: Graphical model representation of the optimal control problem.
We begin by representing the optimal control problem in Eq. (1) as a graphical model. A graphical model depicts dependencies between random variables, where nodes represent random variables and edges represent dependencies. Equation (1) can be represented by the graphical model shown in Fig. 3. As illustrated, $\mathbf{x}_{t+1}$ depends on $\mathbf{u}_{t}$ and $\mathbf{x}_{t}$, representing Eq. (1b). Additionally, edges from $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ lead to $\mathcal{O}_{t}$, called the optimality variable, i.e., a virtual binary variable that returns whether the state and input are optimal ($\mathcal{O}_{t}=1$) or not optimal ($\mathcal{O}_{t}=0$). This $\mathcal{O}_{t}$ represents the cost function in Eq. (1a).
Based on this graphical model, we derive the probability distribution of the optimal trajectory $\tau=[\mathbf{x}_{0:T},\mathbf{u}_{0:T-1}]$. Setting $\mathcal{O}_{0:T}=[\mathcal{O}_{0},\dots,\mathcal{O}_{T}]$, Bayes’ theorem gives the optimal trajectory distribution $p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})$ as:
$p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})$ $\propto\underbrace{p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)}_{\text{Optimality likelihood}}\times\underbrace{p(\tau)}_{\text{Trajectory distribution}}$
$=p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)p(\mathbf{x}_{0})\prod_{t=0}^{T-1}p(\mathbf{x}_{t+1}\mid\mathbf{x}_{t},\mathbf{u}_{t})p(\mathbf{u}_{t}),$ (3)
where $p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)$ is the optimality likelihood, representing the probability that $\tau$ is an optimal trajectory.
3.1.2 Approximating the Optimal Control Distribution via Kullback-Leibler (KL) Divergence Minimization
While we modeled the optimal trajectory distribution $p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})$ in the previous section Eq. (3), the virtual binary variable $\mathcal{O}_{0:T}$ cannot be computed in advance, making this distribution intractable. We therefore employ variational inference to approximate the optimal trajectory distribution. Specifically, we introduce a distribution $\pi(\tau)$ that does not depend on $\mathcal{O}_{0:T}$ and minimize the KL divergence $\mathbb{D}_{\rm{KL}}$ between $p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})$ and $\pi(\tau)$:
$\pi^{*}(\tau)$ $=\min_{\pi}\mathbb{D}_{\rm{KL}}\left(\pi(\tau)\parallel p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})\right)$
$=\min_{\pi}\mathbb{E}_{\pi(\tau)}\left[\log\pi(\tau)-\log p(\tau\mid\mathcal{O}_{0:T}=\mathbf{1})\right]$
$=\min_{\pi}\{\mathbb{E}_{\pi(\tau)}[-\log p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)]$
$\qquad\quad+\mathbb{E}_{\pi(\tau)}\left[\log(\pi(\tau)-\log p(\tau))\right]\},$ (4)
where the final transformation uses Eq. (3). Assuming the optimality likelihood $p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)$ follows a Boltzmann distribution, Eq. (4) can be marginalized over states as:
$\pi^{*}(\mathbf{u}_{0:T-1})$ $=\min_{\pi}\{\lambda^{-1}\overbrace{\mathbb{E}_{\pi(\mathbf{u}_{0:T-1})}\left[J_{\tau}(\mathbf{u}_{0:T-1})\right]}^{\text{Expected cost function}}$
$\;\;\;\;+\underbrace{\mathbb{D}_{\rm{KL}}\left(\pi(\mathbf{u}_{0:T-1})\parallel p(\mathbf{u}_{0:T-1})\right)}_{\text{Regularization term for deviation from prior dist.}}\},$ (5)
where, with normalization constant $\eta$:
$p(\mathcal{O}_{0:T}=\mathbf{1}\mid\tau)=\eta^{-1}\exp\left(-\lambda^{-1}J_{\tau}(\mathbf{u}_{0:T-1})\right),$ (6)
$J_{\tau}(\mathbf{u}_{0:T-1})=\mathbb{E}_{p(\mathbf{x}_{0:T}\mid\mathbf{u}_{0:T-1})}\left[J(\mathbf{x}_{0:T},\mathbf{u}_{0:T-1})\right],$ (7)
are substituted into Eq. (4) [^2] Equation (5) can also be derived via ELBO minimization.. Equation (6) defines the Boltzmann distribution, a probability density function that assigns higher values to trajectories with lower costs. As shown in Fig. 4, the Boltzmann distribution’s shape changes with the temperature parameter $\lambda$: smaller $\lambda$ produces distributions more sharply peaked around the lowest-cost trajectory. While the Boltzmann distribution can be derived from maximum entropy principles, there is no strict theoretical requirement to use it, and alternative distributions are possible [^3] The Cross Entropy Method employs a threshold-based distribution as the optimality likelihood [botev2013cross].. However, the Boltzmann distribution is commonly used due to its practical tractability.
Examining Eq. (5), we see that the optimal control distribution $\pi^{*}$ minimizes a weighted sum (with weight determined by temperature parameter $\lambda$) of the expected cost function and the KL divergence from the prior distribution. Thus, $\pi^{*}(\mathbf{u}_{0:T-1})$ is a distribution that minimizes the expected cost while constraining deviation from the prior distribution.
Figure 4: Cost function and Boltzmann distribution. The Boltzmann distribution assigns higher probability to trajectories with lower costs, with its shape controlled by the temperature parameter $\lambda$. The example cost function is $J(u)=0.6u^{2}\sin(5\pi u)$.
Figure 5: Variation of the optimal control distribution with temperature parameter $\lambda$. The cost function is the same as in Fig. 4. The prior distribution is $p(u)=\mathcal{N}(-2.0,1.0)$.
3.1.3 Computing the Optimal Control Distribution via the Method of Lagrange Multipliers
Finally, we derive Eq. (2) from Eq. (5) using Lagrange multipliers $\alpha$. Specifically, we minimize Eq. (5) subject to the probability normalization constraint $\int\pi(\mathbf{u}_{0:T-1})d\mathbf{u}_{0:T-1}=1$. Defining the Lagrangian as
$\mathcal{L}(\pi,\alpha)$
$=\lambda^{-1}\mathbb{E}_{\pi(\mathbf{u}_{0:T-1})}\left[J_{\tau}(\mathbf{u}_{0:T-1})\right]+\mathbb{D}_{\rm{KL}}\left(\pi(\mathbf{u}_{0:T-1})\parallel p(\mathbf{u}_{0:T-1})\right)+\alpha\left(\int\pi d\mathbf{u}_{0:T-1}-1\right)$
we obtain:
$\frac{\partial\mathcal{L}}{\partial\pi}=\lambda^{-1}J_{\tau}+\log\frac{\pi^{*}}{p}-1+\alpha=0$
$\Leftrightarrow\;\;$ $\pi^{*}=\exp(1-\alpha)\exp\left(-\lambda^{-1}J_{\tau}\right)p(\mathbf{u}_{0:T-1}).$
From the normalization constraint
$\int\pi^{*}d\mathbf{u}_{0:T-1}=\exp(1-\alpha)\int\exp\left(-\lambda^{-1}J_{\tau}\right)p(\mathbf{u}_{0:T-1})d\mathbf{u}_{0:T-1}=1,$
we have
$\exp(1-\alpha)=\left[\int\exp\left(-\lambda^{-1}J_{\tau}\right)p(\mathbf{u}_{0:T-1})d\mathbf{u}_{0:T-1}\right]^{-1}=Z^{-1}.$
Therefore, we finally complete the derivation of Eq. (2).
3.2 Analysis of the Optimal Control Distribution $\pi^{*}$
Having derived the optimal control distribution in Eq. (2) through mathematical manipulations, we now pause to examine its meaning and characteristics.
3.2.1 Influence of the Prior Distribution on the Optimal Control Distribution
As shown in Eq. (2), the optimal control distribution is expressed as the product of the Boltzmann distribution and the control input prior distribution. This means that “optimal actions” are not merely those minimizing the cost function in Eq. (1a), but rather those that also account for the action prior distribution (Fig. 5).
Reference [kappen2005path] illustrates this characteristic with the analogy of a “drunken spider crossing a pond.” Imagine a spider attempting to cross a pond via a narrow bridge. When sober (i.e., when the prior distribution has small variance), the optimal action is to cross the pond via the shortest path using the bridge. However, when drunk (i.e., when the prior distribution has large variance), there is a risk of falling into the pond while crossing the narrow bridge. Therefore, detouring around the pond becomes the optimal action. Thus, optimal actions depend not only on the cost function but also on the prior distribution, which appears to be a principle valid even in nature.
Note that models that use Boltzmann distributions as optimal distributions for sample generation are called Energy-Based Models (EBM), long studied in machine learning [lecun2006tutorial]. The difference between EBM and Eq. (2) lies in the prior distribution, which originates from the dynamics in Eq. (1b).
3.2.2 Decision-Making Through Symmetry Breaking
Figure 6: Example of decision-making through symmetry breaking.
One intriguing phenomenon arising from the prior distribution is “symmetry breaking” [kappen2005path]. Symmetry breaking refers to the phenomenon where the number of modes in the optimal control distribution decreases over time as the noise level of the prior distribution diminishes.
Consider the obstacle avoidance example illustrated in Fig. 6. When far from the obstacle, the optimal control distribution has two modes representing paths around the obstacle. At this point, the probabilities of selecting left or right paths are nearly equal, resulting in no execution of avoidance behavior. However, as time progresses and the robot approaches the obstacle, at some point the prior distribution randomly biases toward one direction. Subsequently, the optimal control distribution, bound to the prior, converges to a single mode over time.
From a control perspective, this mode convergence through symmetry breaking represents delayed decision-making. In deterministic control methods, “early decisions” to proceed left or right occur even far from obstacles. In contrast, probabilistic inference-based methods tend to postpone decisions until reasonably close to obstacles.
3.2.3 Relationship Between Temperature Parameter and Optimal Control Distribution
As shown in Eq. (2), the optimal control distribution depends on the temperature parameter $\lambda$. The temperature parameter originates from Eq. (6) and, as shown in Eq. (5), controls the weight between the expected cost and the regularization term for deviation from the prior.
Figure 5 illustrates how the optimal control distribution changes with the temperature parameter. Larger temperature parameters increase the influence of the prior distribution, making the optimal control distribution smoother and closer to the prior. Conversely, smaller temperature parameters increase the influence of the Boltzmann distribution, producing sharper distributions concentrated near the Boltzmann distribution’s modes. As extreme cases, $\lambda\to 0$ yields deterministic optimal actions that minimize the cost function, while $\lambda\to\infty$ yields stochastic optimal actions following only the prior distribution. Thus, the temperature parameter tunes the balance between deterministic and stochastic characteristics of the optimal control distribution.
3.2.4 Factors Governing Sample Complexity
As we will discuss in Section 3.3, we ultimately approximate the optimal control distribution in Eq. (2) using Monte Carlo methods. Therefore, the optimal control distribution should ideally have a shape that can be approximated with as few samples as possible. The number of samples required to approximate a probability distribution depends on its shape, a property known as sample complexity.
Generally, sample complexity increases as the distribution becomes more peaked, requiring more samples [^4] Intuitively, drawing from a broader “winning region” with random samples requires fewer samples than pinpointing a single “correct answer.”. According to Eq. (2), the main factors determining the optimal control distribution’s shape are:
  • Temperature parameter $\lambda$: As discussed in Section 3.2.3, smaller temperature parameters produce sharper optimal control distributions with higher sample complexity.
  • Prior distribution $p(\mathbf{u}_{0:T-1})$: Sharper priors (i.e., smaller exploration noise) increase sample complexity.
  • Cost function $J_{\tau}$: When the cost function landscape is steep, e.g., when many obstacles lie directly in front of the robot such that many samples collide, sample complexity increases.
Since these interacting factors make estimating required number of samples difficult, in practice, more samples are generally better.
3.3 Control Input Generation from the Optimal Control Distribution: The MPPI Example
The preceding sections focused on how the optimal control distribution is represented and its characteristics. As shown in Fig. 1, we now focus on how to generate control inputs from the optimal control distribution—the essential step for actual control.
Despite having a mathematical expression for the optimal control distribution in Eq. (2), directly sampling from it computationally is challenging. This is because the distribution can be highly complex depending on the cost function and prior, or may be ill-defined until a specific control input sequence is determined. We therefore again employ a variational inference approach. Specifically, we approximate the optimal control distribution $\pi^{*}$ using a simple variational distribution $\pi_{\theta}$:
$\mathbf{u}_{0:T-1}^{*}\sim\pi_{\theta}^{*}=\min_{\theta}\mathbb{D}_{\rm{KL}}\left(\pi^{*}\parallel\pi_{\theta}\right),$ (9)
and find the variational distribution $\pi_{\theta}$ that achieves this minimum. The specific choice of $\pi_{\theta}$ varies by method.
The most widely used probabilistic inference MPC method is MPPI [williams2018information]. MPPI employs a multivariate Gaussian distribution with fixed covariance $\Sigma$ as the variational distribution. Specifically, setting $\theta=\mu_{0:T-1}=[\mu_{0},\dots,\mu_{T-1}]$ and
$\pi_{\mu_{0:T-1}}=\prod_{t=0}^{T-1}\mathcal{N}(\mu_{t},\Sigma),$
$\mu_{0:T-1}$ $=\min_{\mu_{0:T-1}}\mathbb{D}_{\rm{KL}}\left(\pi^{*}\parallel\prod_{t=0}^{T-1}\mathcal{N}(\mu_{t},\Sigma)\right)$
$=\min_{\mu_{0:T-1}}\mathbb{E}_{\pi^{*}}\left[\cancel{\log\pi^{*}}-\log\prod_{t=0}^{T-1}\mathcal{N}(\mu_{t},\Sigma)\right]$
$=\max_{\mu_{0:T-1}}\mathbb{E}_{\pi^{*}}\underbrace{\left[\sum_{t=0}^{T-1}\log\mathcal{N}(\mu_{t},\Sigma)\right]}_{\text{Concave}}$ (10)
$=\max_{\mu_{0:T-1}}\int\pi^{*}\log\mathcal{N}(\mu_{t},\Sigma)d\mathbf{u}_{0:T-1}.$ (11)
Since Eq. (10) is concave, the global optimal solution $\mu_{0:T-1}^{*}$ maximizing Eq. (11) can be found via the stationarity condition. Setting $\mathcal{L}=\int\pi^{*}\log\mathcal{N}(\mu_{t},\Sigma)d\mathbf{u}_{0:T-1}$:
$\frac{\partial\mathcal{L}}{\partial\mu_{t}}$ $=\int\pi^{*}(\mathbf{u}_{0:T-1}-\mu_{0:T-1}^{*})d\mathbf{u}_{0:T-1}=0$
$\Leftrightarrow\;\;\mu_{0:T-1}^{*}$ $=\frac{\int\mathbf{u}_{0:T-1}\exp(-\lambda^{-1}J_{\tau})p(\mathbf{u}_{0:T-1})d\mathbf{u}_{0:T-1}}{\int\exp(-\lambda^{-1}J_{\tau})p(\mathbf{u}_{0:T-1})d\mathbf{u}_{0:T-1}}$
$=\frac{\mathbb{E}_{p(\mathbf{u}_{0:T-1})}\left[\mathbf{u}_{0:T-1}\exp(-\lambda^{-1}J_{\tau}(\mathbf{u}_{0:T-1}))\right]}{\mathbb{E}_{p(\mathbf{u}_{0:T-1})}\left[\exp(-\lambda^{-1}J_{\tau}(\mathbf{u}_{0:T-1}))\right]},$ (12)
where we use Eq. (2). The expectation $\mathbb{E}_{p(\mathbf{u}_{0:T-1})}[\cdot]$ in Eq. (12) can be computed numerically via Monte Carlo methods, with the prior distribution $p(\mathbf{u}_{0:T-1})$ for random sampling typically taken as the optimal solution from the previous time step, i.e.,
$p(\mathbf{u}_{0:T-1})=\mathcal{N}(\mu_{0:T-1}^{\rm{prev}},\Sigma).$
Finally, the MPPI algorithm consists of just three remarkably simple steps:
  1. 1.
    Sample $K$ control inputs from the prior distribution:
    $\mathbf{u}_{0:T-1}^{k}\sim\mathcal{N}(\mu_{0:T-1}^{\rm{prev}},\Sigma)\;\;(k=\{1,\dots,K\})$
  2. 2.
    Forward simulation and calculating trajectory costs using Eq. (7):
    $J_{\tau}^{k}=J_{\tau}(\mathbf{u}_{0:T-1}^{k})\;\;(k=\{1,\dots,K\})$
  3. 3.
    Find an optimal solution via Monte Carlo approximation of Eq. (12):
    $\mu_{0:T-1}^{*}=\sum_{k=1}^{K}\left\{\rm{Softmax}(-\lambda^{-1}J_{\tau}^{k})\mathbf{u}_{0:T-1}^{k}\right\}$ (13)
Note that Eq. (12) can be implemented simply using the Softmax function as shown in Eq. (13), i.e., merely a weighted average of input samples using Softmax weights. Moreover, steps (1) and (2) are independent operations for each sample $k$, enabling straightforward GPU parallelization via PyTorch and similar frameworks [kohonda2025mppi_playground].
A key advantage of MPPI is that it finds the global optimal solution (in the sense of KL divergence minimization) through a single round of sampling and weighted averaging [^5] However, as discussed in Section 3.2.4, the required number of samples depends on the configuration.. This stems from the concavity of Eq. (10), which results from fixing the variance of the multivariate Gaussian variational distribution. If we simultaneously optimize the variance, Eq. (11) would no longer be concave, requiring iterative repetition of the three steps above in a single optimization, increasing computational cost. However, since the variational distribution’s variance setting significantly impacts performance, careful configuration is necessary [honda2024stein].
4 Applications and Extensions
The previous sections presented the theoretical framework of probabilistic inference-based MPC. The general scheme can be summarized by the following equation:
$\underset{\text{Variational dist.}}{\pi^{*}_{\theta}(\mathbf{u}_{0:T-1})}\mathrel{\overset{\text{min.\ KL div.}}{\longleftrightarrow}}\underset{\text{Optimal control dist.}}{\pi^{*}(\mathbf{u}_{0:T-1})}\propto\underset{\text{Boltzmann dist.}}{\exp\left(-\lambda^{-1}J_{\tau}(\mathbf{u}_{0:T-1})\right)}\;\times\;\underset{\text{Prior dist.}}{p(\mathbf{u}_{0:T-1})}.$ (14)
This equation indicates that the optimal control distribution is expressed as the product of the Boltzmann distribution and the prior distribution (Eq. (2)), and that the variational distribution approximates this optimal control distribution by minimizing the KL divergence. We can generate control input sequences by sampling from the variational distribution via Eq. (9). This section discusses key points for improving performance and tuning, as well as advanced research topics based on this general framework.
4.1 Keys to Performance Improvement and Tuning
Revisiting Eq. (14), for a given specific task and robot, the trajectory cost $J_{\tau}$ and the control input $\mathbf{u}_{0:T-1}$ are uniquely determined. Additionally, the optimal control distribution is theoretically given by Eq. (2), although there have been discussions regarding the shape of the optimality likelihood.
Therefore, the design elements we need to consider are the temperature parameter $\lambda$, the prior distribution $p(\mathbf{u}{0:T-1})$, and the variational distribution $\pi{\theta}$. Among these, as discussed in Section 3.2.3, the temperature parameter is a hyperparameter that governs the trade-off between stochastic and deterministic behavior. In contrast, the design of the prior and variational distributions has a significant influence on control performance and characteristics. Based on this insight, we discuss these two design aspects, i.e., the prior and variational distributions, in detail below.
4.1.1 Prior Distribution $p(\mathbf{u}_{0:T-1})$ is “Search Region for Solutions”
Figure 7: Variation of the optimal control distribution with the prior distribution. The cost function is the same as in Fig. 4. Four cases of the prior distribution $p(u)=\mathcal{N}(\mu,\sigma)$ are shown with $\mu=\{-3.0,0.0\}$ and $\sigma=\{0.25,1.0\}$.
As discussed in Section 3.2.1, the optimal control distribution is regularized by the prior distribution, making the prior a crucial element determining the optimal control distribution’s shape.
Figure 7 shows how the optimal control distribution changes with variations in the prior distribution’s mean and variance. Increasing the deviation between the prior distribution and the Boltzmann distribution’s principal mode causes the optimal control distribution to deviate from the Boltzmann distribution, degrading optimality in the sense of cost function minimization (bottom-right and top-right panels). Additionally, increasing the prior distribution’s variance moves the optimal control distribution closer to the Boltzmann distribution when their deviation is large (top-right), which is desirable. However, when their deviation is small (top-left), the resulting optimal control distribution also has large variance, lowering the probability of generating the optimal control distribution’s principal mode and potentially failing to achieve optimal actions in the cost function sense. Thus, the prior distribution to avoid most is “one with large deviation from the Boltzmann distribution and small variance” (bottom-right), while the most desirable prior is “one close to the Boltzmann distribution with small variance” (bottom-left).
This becomes more intuitive from an algorithmic perspective. As shown in Eq. (12), the prior distribution defines the “search region for solutions” from which we actually sample. Therefore, it is desirable to focus sampling around the vicinity of the Boltzmann distribution’s mode peaks.
In summary, the prior distribution must be designed with awareness of the Boltzmann distribution (i.e., the cost function). Consequently, various approaches have been proposed, including cost-aware adaptive priors [honda2024stein, yi2024covo], optimization via learning [power2022variational], and utilization of reinforcement learning policies [hansen2022temporal].
4.1.2 Variational Distribution $\pi_{\theta}$ is “Diversity of Behaviors”
The variational distribution approximates the optimal control distribution and ultimately generates control inputs. Therefore, the variational distribution determines the “diversity of generated actions”. For example, MPPI assumes a unimodal Gaussian variational distribution, which cannot capture multiple modes of the behavior, i.e., a multimodal optimal control distribution, generating solutions only near a single mode.
To improve behavior diversity, more flexible variational distributions are needed. Consequently, Gaussian mixture distributions [okada2020variational] and particle-based representations [lambert2021stein] have been proposed as variational distributions. These methods enhance exploration capabilities through increased solution diversity, facilitating model-based reinforcement learning [okada2020variational]. However, they tend to suffer from increased computational cost and reduced stability.
4.2 Advanced Research Topics
4.2.1 Theoretical Analysis
Probabilistic inference MPC is even more challenging to analyze theoretically than numerical optimization-based MPC. However, recent years have seen a small number of theoretical analysis studies. Reference [yi2024covo] proves that MPPI guarantees linear convergence when the number of samples is infinite and the cost function is quadratic. However, infinite samples are impractical. It is important to note that with finite samples, convergence is not guaranteed even for quadratic cost functions due to sample complexity effects (Section 3.2.4).
4.2.2 Accelerating Decision-Making
As discussed in Section 3.2.2, optimal control distributions in probabilistic inference MPC tend to exhibit delayed decision-making. Additionally, the variational distribution obtained in Eq. (9) minimizes the forward KL divergence. Forward KL divergence minimization corresponds to maximum likelihood estimation, causing the resulting variational distribution to tend to cover multiple modes. Consequently, decision-making is further delayed.
This decision-making delay can sometimes lead to dangerous behaviors. For instance, when controlling fast-moving vehicles, optimal behavior involves avoiding obstacles with minimal clearance, so even slight dynamics errors may cause collisions. To accelerate decision-making, methods that minimize reverse KL divergence instead of forward KL divergence in Eq. (9) have been proposed [honda2024stein, kobayashi2022real]. Reverse KL divergence minimization tends to cover only a single mode rather than multiple modes, enabling faster decision-making.
4.2.3 Connection to Diffusion Models
Recently, diffusion models that generate diverse data through denoising have attracted attention, and connections between MPPI and diffusion models have been identified [xue2024full]. Literature [xue2024full] demonstrates that when solving the optimal control distribution using score-based diffusion models, part of the score function can be replaced with MPPI computation, enabling sample generation by solving MPPI for the number of diffusion steps without requiring training. This method can capture smaller modes compared to MPPI and has been reported to achieve high performance even in high-dimensional control problems such as quadruped robots. However, the need to solve MPPI for each diffusion step results in very high computational cost per sample generation, which remains a challenge for future work.
5 Conclusion
Probabilistic inference-based MPC is a framework that combines theoretical depth with implementation simplicity. We anticipate it will serve as a foundational tool not only in robotics but across diverse fields in the future. We hope this article inspires readers in their research across various domains and contributes to further developments in the field.
[…]